OpenStudy (anonymous):

A country's population in 1992 was 222 million. In 2001 it was 224 million. Estimate the population in 2004 using the exponential growth formula. Round your answer to the nearest million. P = Aekt

OpenStudy (anonymous):

@Pizza7

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@aaronq

OpenStudy (amistre64):

i believe kt is an exponent here?

OpenStudy (anonymous):

yes @amistre64

OpenStudy (amistre64):

id just use the 2 points to define the parameters with

OpenStudy (amistre64):

1992 was 222 million. t=0, P=222 2001 it was 224 million t=9, P=224 \[222=Ae^{0k}\] \[224=Ae^{9k}\] well, A has to be 222 sooo \[224=222e^{9k}\] log it out to find k and have a descent formula

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

no... Im confused

OpenStudy (amistre64):

youll have to tell me the confusion then, since this is too simple for me to dissect any further to me.

OpenStudy (anonymous):

What do you mean to log it out?

OpenStudy (amistre64):

we have to solve for k to find the rate at which the population is changing. e^k has an inverse. its the natural log function

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

this is the process i have in mind when i say log it out. \[P=Ae^{kt}\] \[P/A=e^{kt}\] \[ln(P/A)=ln(e^{kt})\] \[ln(P/A)=kt~ln(e)\] \[ln(P/A)=kt\] \[\frac{ln(P/A)}{t}=k\]

OpenStudy (anonymous):

OH! ok! :)

OpenStudy (amistre64):

:) once we know k, then its just a matter of t=12 i believe \[P=222e^{12k}\]is our approximation

OpenStudy (anonymous):

Thank you! :D

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

k=0.00074738916 Or 0.0007